OpenStudy (anonymous):

What is the simplified form of 4z^2-16z+15/2z^2+z-15

OpenStudy (pfenn1):

\[\frac{4z^2-16z+15}{2z^2+z-15}\]Have you tried factoring the numerator and denominator?

OpenStudy (anonymous):

idk how

OpenStudy (anonymous):

i think its 2z-3/x-3?

OpenStudy (anonymous):

is that right?

OpenStudy (pfenn1):

\[\frac{(2z-3)\cancel {(2z-5)}}{\cancel {(2z-5)}(z+3)}=\frac{2z-3}{z+3}\]This is what I got. Sorry it took so long. I was getting dinner ready.